OpenStudy (chris215):

Find the exact value of the following limit: (e^(6x)-6x-1)/(x^2)

OpenStudy (chris215):

I got 18

OpenStudy (solomonzelman):

x --> what ?

OpenStudy (chris215):

as x->0

OpenStudy (mathmale):

By (e^(6x)-6x-1)/(x^2) do you mean the following?\[\frac{ e ^{6x}-6x-1 }{ x^2 }?\]

OpenStudy (solomonzelman):

\(\large\color{#000000}{\displaystyle\lim_{x \rightarrow ~0} \frac{e^{6x}-6x-1}{x^2} }\)

OpenStudy (chris215):

yes

OpenStudy (solomonzelman):

use L'Hospital's rule. (it is 0/0 now)

OpenStudy (mathmale):

I agree with Solomon: Use l'H's Rule. Differentiate the numerator with respect to x. Differentiate the denominator with respect to x, separately. Form a fraction out of your results. Take the limit as x -> 0 of that fraction.

OpenStudy (solomonzelman):

differentiate, and if you have questions, ask.

OpenStudy (mathmale):

@chris215 : Share your thoughts, please, and especially share any work you've been able to do on this problem. Have you worked with l;Hopital's Rule before?

OpenStudy (solomonzelman):

If you used L'Hospital's rule twice, then 18 is right

OpenStudy (solomonzelman):

well, 18 is just correct...

OpenStudy (chris215):

I differentiated and got d/dx(e^(6 x)-6 x-1) = 6 (e^(6 x)-1)

OpenStudy (chris215):

and x^2= 2x

OpenStudy (solomonzelman):

yes, very good \(\large\color{#000000}{\displaystyle\lim_{x \rightarrow ~0} \frac{e^{6x}-6x-1}{x^2} }\) since it is not 0/0 when x=0 is plugged, we differentiate top and bottom. \(\large\color{#000000}{\displaystyle\lim_{x \rightarrow ~0} \frac{\frac{d}{dx}\left[e^{6x}-6x-1\right] }{\frac{d}{dx}\left[x^2\right]} }\) \(\large\color{#000000}{\displaystyle\lim_{x \rightarrow ~0} \frac{6e^{6x}-6}{2x} }\) \(\large\color{#000000}{\displaystyle\lim_{x \rightarrow ~0} \frac{3(e^{6x}-1)}{x} }\) 0/0 again, and thus differentiate again...

OpenStudy (solomonzelman):

you will use the chain rule for the exponent again, and then evaluate the limit.... and you will get 18 - the correct result you already had.

OpenStudy (solomonzelman):

((just asking, when you evaluated the limit first, you used an alternative approach... was it graphing? ))

OpenStudy (zarkon):

taylor series is the quickest way

OpenStudy (mathmale):

Please give @chris215 an opportunity to respond.

OpenStudy (solomonzelman):

Zarkon, honestly, that is the first time I have ever heard this from anyone. To me (and to many others), finding L'H'S twice and differentiating is a matter of 1 minute...

OpenStudy (chris215):

yeah I used graphinng

OpenStudy (solomonzelman):

but, whatever you say :)

OpenStudy (solomonzelman):

Yeah, graphing is good, but not always available (eg tests)

OpenStudy (chris215):

true but thanks so much! that wa very helpful!

OpenStudy (solomonzelman):

Anytime:) Just remember that if you see that when you plug in the variable into the limit you get 0/0 or ∞/∞ then you can differentiate top and bottom (this is L'Hospital's rule)....